# Utilizing and extending superconducting circuit toolbox to simulate analog quantum gravity

### v2

@@ Appendix A Josephson junction array with gyrators
$\mathcal{H}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L}\left(\phi_{j+1}-\phi_{j}\right)^{2}+M\phi_{j}^{2}\ ,$ (43)
@@
$\mathcal{H}=E_{C}\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\left[N_{m}N_{-m}+i2G\sin\left(\frac{2\pi m}{J}\right)\left\{N_{m},\phi_{m}\right\}+\beta_{m}\phi_{m}\phi_{-m}\right]\ ,$ (45)
@@
$\beta_{m}$ $=\frac{4E_{C}G^{2}\sin^{2}\left(\frac{2\pi m}{J}\right)+4E_{L}\sin^{2}\left(\frac{\pi m}{J}\right)+M}{E_{C}}.$
Now, we write the Fourier transformed charge and phase operators in terms of bosonic annihilation and creation operators
$N_{m}$ $=\frac{i}{\sqrt{2}}\left(a_{-m}-a_{m}^{\dagger}\right)\ ,$ (46)
$\phi_{m}$ $=\frac{1}{\sqrt{2}}\left(a_{m}+a_{-m}^{\dagger}\right)\ ,$ (47)
that obey the following commutation relations $[a_{m},a^{\dagger}_{m^{\prime}}]=\delta_{mm^{\prime}}$, $[a_{m},a_{m^{\prime}}]=0$. Finally, the Hamiltonian becomes
$\mathcal{H}$ $=\frac{E_{C}}{2}\sum_{m=1}^{\frac{J}{2}-1}\left(\begin{array}[]{cccc}a_{m}^{\dagger}&a_{-m}^{\dagger}&a_{m}&a_{-m}\end{array}\right)H_{m}\left(\begin{array}[]{c}a_{m}\\ a_{-m}\\ a_{m}^{\dagger}\\ a_{-m}^{\dagger}\end{array}\right)+\frac{E_{C}}{2}\left(\begin{array}[]{cc}a_{0}^{\dagger}&a_{0}\end{array}\right)h\left(\begin{array}[]{c}a_{0}\\ a_{0}^{\dagger}\end{array}\right)+\frac{E_{C}}{2}\left(\begin{array}[]{cc}a_{\frac{J}{2}}^{\dagger}&a_{\frac{J}{2}}\end{array}\right)h^{\prime}\left(\begin{array}[]{c}a_{\frac{J}{2}}\\ a_{\frac{J}{2}}^{\dagger}\end{array}\right)\ ,$ (59)
$\mathcal{H}$ $=\sum_{m=1}^{\frac{J}{2}-1}\mathcal{H}_{1,m}+\mathcal{H}_{2}+\mathcal{H}_{3}\ ,$ (60)
@@
$H_{m}$ $=\left(\beta_{m}+1\right)\mathbb{I}_{4}+4G\sin\left(\frac{2\pi m}{J}\right)\left(\begin{array}[]{cc}\sigma_{z}&0\\ 0&\sigma_{z}\end{array}\right)+\left(\beta_{m}-1\right)\left(\begin{array}[]{cc}0&\sigma_{x}\\ \sigma_{x}&0\end{array}\right)\ ,$ (65)
$h$ $=\left(\beta_{0}+1\right)\mathbb{I}_{2}+\left(\beta_{0}-1\right)\sigma_{x}\ ,$ (66)
$h^{\prime}$ $=\left(\beta_{\frac{J}{2}}+1\right)\mathbb{I}_{2}+\left(\beta_{\frac{J}{2}}-1\right)\sigma_{x}.$ (67)
$\mathbb{I}_{n}$ is a $n$-dimensional identity matrix while $\sigma_{z}$ and $\sigma_{x}$ are Pauli matrices. We will focus on diagonalization of $\mathcal{H}_{1,m}$ in eq.(60), the same formalism can be used for the other parts.
@@
$\mathcal{H}_{1,m}$ $=\frac{E_{C}}{2}\left(\begin{array}[]{cccc}b_{m}^{\dagger}&b_{-m}^{\dagger}&b_{m}&b_{-m}\end{array}\right)\tau_{z}\Lambda_{m}\left(\begin{array}[]{c}b_{m}\\ b_{-m}\\ b_{m}^{\dagger}\\ b_{-m}^{\dagger}\end{array}\right)\ ,$ (73)
@@
$\mathcal{H}_{1,m}$ $=\frac{E_{C}}{2}\left(\begin{array}[]{cccc}a_{m}^{\dagger}&a_{-m}^{\dagger}&a_{m}&a_{-m}\end{array}\right)\tau_{z}Q_{m}\tau_{z}\tau_{z}\Lambda_{m}Q_{m}^{-1}\left(\begin{array}[]{c}a_{m}\\ a_{-m}\\ a_{m}^{\dagger}\\ a_{-m}^{\dagger}\end{array}\right)$ (83)
$=\frac{E_{C}}{2}\left(\begin{array}[]{cccc}b_{m}^{\dagger}&b_{-m}^{\dagger}&b_{m}&b_{-m}\end{array}\right)\tau_{z}\Lambda_{m}\left(\begin{array}[]{c}b_{m}\\ b_{-m}\\ b_{m}^{\dagger}\\ b_{-m}^{\dagger}\end{array}\right).$ (89)
@@
$\mathcal{H}$ $=\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\omega_{m}b_{m}^{\dagger}b_{m},$ (90)
$\omega_{m}$ $=2\sqrt{E_{C}^{2}\beta_{m}}+4E_{C}G\sin\left(\frac{2\pi m}{J}\right).$ (91)
@@
$\mathcal{H}^{\prime}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L}\left(\phi_{j+1}-\phi_{j-1}\right)^{2}+M\phi_{j}^{2}.$ (92)
@@
$\mathcal{H}^{\prime}$ $=\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\omega_{m}b_{m}^{\dagger}b_{m},$ (93)
$\omega_{m}$ $=2\sqrt{E_{C}^{2}\beta_{m}}+4E_{C}G\sin\left(\frac{2\pi m}{J}\right),$ (94)
@@ Appendix B Bogoliubov transformation for calculating correlations after quench
$\mathcal{H}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L,j}\left(\phi_{j+1}-\phi_{j}\right)^{2}+M\phi_{j}^{2}.$ (95)
@@
$N_{j}$ $=\frac{i}{\sqrt{2}}\left(a_{j}-a_{j}^{\dagger}\right),$ (96)
$\phi_{j}$ $=\frac{1}{\sqrt{2}}\left(a_{j}+a_{j}^{\dagger}\right),$ (97)
that obey the following commutation relations $[a_{j},a^{\dagger}_{j^{\prime}}]=\delta_{jj^{\prime}}$, $[a_{j},a_{j^{\prime}}]=0$. This allows us to write the Hamiltonian as
$\mathcal{H}=\mathbf{a}^{\dagger}H\mathbf{a},$ (98)
@@
$=\mathbf{b}^{\dagger}Z\Lambda_{<}\mathbf{b}$ (99)
$=\sum_{j=1}^{J}\left(\lambda_{<,j}b^{\dagger}_{j}b_{j}+\lambda_{<,j}b_{j}b^{\dagger}_{j}\right).$ (100)
@@
and operators $\mathbf{b}^{\left(\dagger\right)}$ have the same commutation relations as $\mathbf{a}^{\left(\dagger\right)}$. The matrix $Q$ satisfies the conditions in eq.(75) and (76), while the diagonal matrix $\Lambda_{<}$ satisfies the eq.(77). In position space the Hamiltonian does not break down into independent blocks of $4*4$ matrices, therefore the size of the matrices in the above equations have to be changed appropriately.
@@
$E_{L,j}=\begin{cases}-E_{L}&j_{0}<j<j_{1},\\ E_{L}&\text{everywhere else,}\end{cases}$ (101)
@@
$=\mathbf{a}^{\dagger}ZPZZ\Lambda_{>}P^{-1}\mathbf{a}.$ (102)
The diagonal matrix $\Lambda_{>}$ is not necessarily real, therefore the eq.(75) and (76) doesn’t hold, but these conditions translate to conditions on the first quantized matrix as
$H_{>}$ $=H_{>}^{\dagger},$ (103)
$H_{>}^{T}$ $=XH_{>}X,$ (104)
$X$ $=\left(\begin{array}[]{cc}0&\mathbb{I}_{J}\\ -\mathbb{I}_{J}&0\end{array}\right).$ (107)
@@
$P^{-1}$ $=ZXP^{T}XZ,$ (108)
$\Lambda_{>}$ $=-X\Lambda_{>}X,$ (109)
$\Lambda_{>}$ $=M^{-1}\Lambda^{*}_{>}M,$ (110)
where $M=ZP^{\dagger}ZP$. One might notice that eq.(108) is just the combination of eq.(75) and (76). The unquenched Hamiltonian can now be written as
$\mathcal{H}_{>}$ $=\mathbf{c}^{\bullet}Z_{2J}\Lambda_{>}\mathbf{c}$ (111)
$=\sum_{j=1}^{J}\left(\lambda_{>,j}c^{\bullet}_{j}c_{j}+\lambda_{>,j}c_{j}c^{\bullet}_{j}\right),$ (112)
@@
The new operators obey the following commutation relations $[c_{j},c^{\bullet}_{j^{\prime}}]=\delta_{jj^{\prime}}$, $[c_{j},c_{j^{\prime}}]=0$, $[c^{\bullet}_{j},c^{\bullet}_{j^{\prime}}]=0$. These relations closely resemble Bosonic commutation relations, but it is important to note that the operators $c^{\bullet}_{j}$ and $c_{j}$ are not related by Hermitian conjugation.
@@
$F(t,t^{\prime})\equiv\left<\mathbf{a}(t)\mathbf{a}^{\dagger}(t^{\prime})\right>,$ (113)
where the average is taken over the ground state of the equilibrium system. For a total number of $J$ sites $F$ is a $2J*2J$ matrix with the substructure form
$F(t,t^{\prime})=\left(\begin{array}[]{cc}A(t,t^{\prime})&B(t,t^{\prime})\\ C(t,t^{\prime})&D(t,t^{\prime})\end{array}\right),$ (116)
where $A,B,C\text{ and }D$ are block matrices of dimensions $J*J$ that contain the correlations of the form $\left\langle a_{j}(t)a_{j^{\prime}}^{\dagger}(t^{\prime})\right\rangle$ ,$\left\langle a_{j}^{\dagger}(t)a_{j^{\prime}}^{\dagger}(t^{\prime})\right\rangle$ ,$\left\langle a_{j}(t)a_{j^{\prime}}(t^{\prime})\right\rangle$ and $\left\langle a_{j}^{\dagger}(t)a_{j^{\prime}}(t^{\prime})\right\rangle$ respectively.
@@
$F_{<}(t,t^{\prime})=Q\left<\mathbf{b}(t)\mathbf{b}^{\dagger}(t^{\prime})\right>ZQ^{-1}Z.$ (117)
@@
$F_{<}(t,t^{\prime})=Qe^{-i2(t-t^{\prime})\Lambda_{<}}\frac{\mathbb{I}_{2J}+Z}{2}Q^{-1}Z.$ (118)
@@
$F_{>}(t,t^{\prime})$ =$\mathbf{a}(t)\mathbf{a}^{\dagger}(t^{\prime})\right>$
@@
$=Pe^{-i2\Lambda_{>}t}P^{-1}Q\frac{\mathbb{I}_{2J}+Z}{2}Q^{-1}Pe^{i2\Lambda_{>}t^{\prime}}P^{-1}Z.$ (119)

### v4

@@ Appendix A Josephson junction array with gyrators
$\mathcal{H}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L}\left(\phi_{j+1}-\phi_{j}\right)^{2}+M\phi_{j}^{2}\ ,$ (A.1)
@@
$\mathcal{H}=E_{C}\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\left[N_{m}N_{-m}+i2G\sin\left(\frac{2\pi m}{J}\right)\left\{N_{m},\phi_{m}\right\}+\beta_{m}\phi_{m}\phi_{-m}\right]\ ,$ (A.3)
@@
and $\{\cdot,\cdot\}$ denotes the anti-commutator. Now, we write the Fourier transformed charge and phase operators in terms of bosonic annihilation and creation operators
$N_{m}$ $=\frac{i}{\sqrt{2}}\left(a_{-m}-a_{m}^{\dagger}\right)\ ,$ (A.4)
$\phi_{m}$ $=\frac{1}{\sqrt{2}}\left(a_{m}+a_{-m}^{\dagger}\right)\ ,$ (A.5)
that obey the following commutation relations $a_{m},a^{\dagger}_{m^{\prime}}]=\delta_{mm^{\prime}}$, $a_{m},a_{m^{\prime}}]=0$. Finally, the Hamiltonian becomes
$\mathcal{H}$ $=\frac{E_{C}}{2}\sum_{m=1}^{\frac{J}{2}-1}\left(\begin{array}[]{cccc}a_{m}^{\dagger}&a_{-m}^{\dagger}&a_{m}&a_{-m}\end{array}\right)H_{m}\left(\begin{array}[]{c}a_{m}\\ a_{-m}\\ a_{m}^{\dagger}\\ a_{-m}^{\dagger}\end{array}\right)$ (A.11)
$\quad+\frac{E_{C}}{2}\left(\begin{array}[]{cc}a_{0}^{\dagger}&a_{0}\end{array}\right)h\left(\begin{array}[]{c}a_{0}\\ a_{0}^{\dagger}\end{array}\right)+\frac{E_{C}}{2}\left(\begin{array}[]{cc}a_{\frac{J}{2}}^{\dagger}&a_{\frac{J}{2}}\end{array}\right)h^{\prime}\left(\begin{array}[]{c}a_{\frac{J}{2}}\\ a_{\frac{J}{2}}^{\dagger}\end{array}\right)\ ,$ (A.18)
$\mathcal{H}$ $=\sum_{m=1}^{\frac{J}{2}-1}\mathcal{H}_{1,m}+\mathcal{H}_{2}+\mathcal{H}_{3}\ ,$ (A.19)
@@
$H_{m}$ $=\left(\beta_{m}+1\right)\mathbb{I}_{4}+4G\sin\left(\frac{2\pi m}{J}\right)\left(\begin{array}[]{cc}\sigma_{z}&0\\ 0&\sigma_{z}\end{array}\right)+\left(\beta_{m}-1\right)\left(\begin{array}[]{cc}0&\sigma_{x}\\ \sigma_{x}&0\end{array}\right)\ ,$ (A.24)
$h$ $=\left(\beta_{0}+1\right)\mathbb{I}_{2}+\left(\beta_{0}-1\right)\sigma_{x}\ ,$ (A.25)
$h^{\prime}$ $=\left(\beta_{\frac{J}{2}}+1\right)\mathbb{I}_{2}+\left(\beta_{\frac{J}{2}}-1\right)\sigma_{x}.$ (A.26)
$\mathbb{I}_{n}$ is a $n$-dimensional identity matrix while $\sigma_{z}$ and $\sigma_{x}$ are Pauli matrices. We will focus on diagonalization of $\mathcal{H}_{1,m}$ in eq.(A.19), the same formalism can be used for the other parts.
@@
$\mathcal{H}_{1,m}$ $=\frac{E_{C}}{2}\left(\begin{array}[]{cccc}b_{m}^{\dagger}&b_{-m}^{\dagger}&b_{m}&b_{-m}\end{array}\right)\tau_{z}\Lambda_{m}\left(\begin{array}[]{c}b_{m}\\ b_{-m}\\ b_{m}^{\dagger}\\ b_{-m}^{\dagger}\end{array}\right)\ ,$ (A.32)
@@
$\mathcal{H}_{1,m}$ $=\frac{E_{C}}{2}\left(\begin{array}[]{cccc}a_{m}^{\dagger}&a_{-m}^{\dagger}&a_{m}&a_{-m}\end{array}\right)\tau_{z}Q_{m}\tau_{z}\tau_{z}\Lambda_{m}Q_{m}^{-1}\left(\begin{array}[]{c}a_{m}\\ a_{-m}\\ a_{m}^{\dagger}\\ a_{-m}^{\dagger}\end{array}\right)$ (A.42)
$=\frac{E_{C}}{2}\left(\begin{array}[]{cccc}b_{m}^{\dagger}&b_{-m}^{\dagger}&b_{m}&b_{-m}\end{array}\right)\tau_{z}\Lambda_{m}\left(\begin{array}[]{c}b_{m}\\ b_{-m}\\ b_{m}^{\dagger}\\ b_{-m}^{\dagger}\end{array}\right).$ (A.48)
@@
$\mathcal{H}$ $=\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\omega_{m}b_{m}^{\dagger}b_{m},$ (A.49)
$\omega_{m}$ $=2\sqrt{E_{C}^{2}\beta_{m}}+4E_{C}G\sin\left(\frac{2\pi m}{J}\right).$ (A.50)
@@
$\mathcal{H}^{\prime}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L}\left(\phi_{j+1}-\phi_{j-1}\right)^{2}+M\phi_{j}^{2}.$ (A.51)
@@
$\mathcal{H}^{\prime}$ $=\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\omega_{m}b_{m}^{\dagger}b_{m},$ (A.52)
$\omega_{m}$ $=2\sqrt{E_{C}^{2}\beta_{m}}+4E_{C}G\sin\left(\frac{2\pi m}{J}\right),$ (A.53)
@@ Appendix B Bogoliubov transformation for calculating correlations after quench
$\mathcal{H}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L,j}\left(\phi_{j+1}-\phi_{j}\right)^{2}+M\phi_{j}^{2}.$ (B.1)
@@
$N_{j}$ $=\frac{i}{\sqrt{2}}\left(a_{j}-a_{j}^{\dagger}\right),$ (B.2)
$\phi_{j}$ $=\frac{1}{\sqrt{2}}\left(a_{j}+a_{j}^{\dagger}\right),$ (B.3)
that obey the following commutation relations $a_{j},a^{\dagger}_{j^{\prime}}]=\delta_{jj^{\prime}}$, $a_{j},a_{j^{\prime}}]=0$. This allows us to write the Hamiltonian as
$\mathcal{H}=\mathbf{a}^{\dagger}H\mathbf{a},$ (B.4)
@@
$=\mathbf{b}^{\dagger}Z\Lambda_{<}\mathbf{b}$ (B.5)
$=\sum_{j=1}^{J}\left(\lambda_{<,j}b^{\dagger}_{j}b_{j}+\lambda_{<,j}b_{j}b^{\dagger}_{j}\right).$ (B.6)
@@
and operators $\mathbf{b}^{\left(\dagger\right)}$ have the same commutation relations as $\mathbf{a}^{\left(\dagger\right)}$. The matrix $Q$ satisfies the conditions in Eq. (A.34) and (A.35), while the diagonal matrix $\Lambda_{<}$ satisfies the Eq. (A.36). In position space the Hamiltonian does not break down into independent blocks of 444*4 matrices, therefore the size of the matrices in the above equations have to be changed appropriately.
@@
$E_{L,j}=\begin{cases}-E_{L}&j_{0}<j<j_{1},\\ E_{L}&\text{everywhere else,}\end{cases}$ (B.7)
@@
$=\mathbf{a}^{\dagger}ZPZZ\Lambda_{>}P^{-1}\mathbf{a}.$ (B.8)
The diagonal matrix $\Lambda_{>}$ is not necessarily real, therefore the Eq. (A.34) and (A.35) doesn’t hold, but these conditions translate to conditions on the first quantized matrix as
$H_{>}$ $=H_{>}^{\dagger},$ (B.9)
$H_{>}^{T}$ $=XH_{>}X,$ (B.10)
$X$ $=\left(\begin{array}[]{cc}0&\mathbb{I}_{J}\\ -\mathbb{I}_{J}&0\end{array}\right).$ (B.13)
@@
$P^{-1}$ $=ZXP^{T}XZ,$ (B.14)
$\Lambda_{>}$ $=-X\Lambda_{>}X,$ (B.15)
$\Lambda_{>}$ $=M^{-1}\Lambda^{*}_{>}M,$ (B.16)
where $M=ZP^{\dagger}ZP$. One might notice that Eq. (B.14) is just the combination of Eq. (A.34) and (A.35). The unquenched Hamiltonian can now be written as
$\mathcal{H}_{>}$ $=\mathbf{c}^{\bullet}Z_{2J}\Lambda_{>}\mathbf{c}$ (B.17)
$=\sum_{j=1}^{J}\left(\lambda_{>,j}c^{\bullet}_{j}c_{j}+\lambda_{>,j}c_{j}c^{\bullet}_{j}\right),$ (B.18)
@@
The new operators obey the following commutation relations $c_{j},c^{\bullet}_{j^{\prime}}]=\delta_{jj^{\prime}}$, $c_{j},c_{j^{\prime}}]=0$, $c^{\bullet}_{j},c^{\bullet}_{j^{\prime}}]=0$. These relations closely resemble Bosonic commutation relations, but it is important to note that the operators $c^{\bullet}_{j}$ and $c_{j}$ are not related by Hermitian conjugation.
@@
$F(t,t^{\prime})\equiv\left<\mathbf{a}(t)\mathbf{a}^{\dagger}(t^{\prime})\right$>, (B.19)
where the average is taken over the ground state of the equilibrium system. For a total number of $J$ sites $F$ is a 2J2J2J*2J matrix with the substructure form
$F(t,t^{\prime})=\left(\begin{array}[]{cc}A(t,t^{\prime})&B(t,t^{\prime})\\ C(t,t^{\prime})&D(t,t^{\prime})\end{array}\right),$ (B.22)
where $A,B,C\text{ and }D$ are block matrices of dimensions JJJ*J that contain the correlations of the form $\left\langle a_{j}(t)a_{j^{\prime}}^{\dagger}(t^{\prime})\right\rangle$ $\left\langle a_{j}^{\dagger}(t)a_{j^{\prime}}^{\dagger}(t^{\prime})\right\rangle$ $\left\langle a_{j}(t)a_{j^{\prime}}(t^{\prime})\right\rangle$ and $\left\langle a_{j}^{\dagger}(t)a_{j^{\prime}}(t^{\prime})\right\rangle$ respectively.
@@
$F_{<}(t,t^{\prime})=Q\left<\mathbf{b}(t)\mathbf{b}^{\dagger}(t^{\prime})\right>ZQ^{-1}Z.$ (B.23)
@@
$F_{<}(t,t^{\prime})=Qe^{-i2(t-t^{\prime})\Lambda_{<}}\frac{\mathbb{I}_{2J}+Z}{2}Q^{-1}Z.$ (B.24)
@@
$F_{>}(t,t^{\prime})$ $=\left<\mathbf{a}(t)\mathbf{a}^{\dagger}(t^{\prime})\right$>
@@
$=Pe^{-i2\Lambda_{>}t}P^{-1}Q\frac{\mathbb{I}_{2J}+Z}{2}Q^{-1}Pe^{i2\Lambda_{>}t^{\prime}}P^{-1}Z$
$=e^{-i2H_{>}t}Q\frac{\mathbb{I}_{2J}+Z}{2}Q^{-1}e^{i2H_{>}t^{\prime}}Z.$ (B.25)